OpenStudy (anonymous):

what are the next two numbers in the pattern bellow 7,21,63,189..... Medals Will Be Given

OpenStudy (john_es):

Simply, multiply by three, 3*189= 3*3*189= 3*3*3*189=

OpenStudy (anonymous):

7,21,63,189,567,1701 @ shankar1